### v2

# Gesture Recognition from body-Worn RFID under Missing Data

Abstract
We explore hand-gesture recognition through the use of passive body-worn reflective tags. A data processing pipeline is proposed to address the issue of missing data. Specifically, missing information is recovered through linear and exponential interpolation and extrapolation. Furthermore, imputation and proximity-based inference are employed. We represent tags as nodes in a temporal graph, with edges formed based on correlations between received signal strength (RSS) and phase values across successive timestamps, and we train a graph-based convolutional neural network that exploits graph-based self-attention. The system outperforms state-of-the-art methods with an accuracy of 98.13% for the recognition of 21 gestures. We achieve 89.28% accuracy under leave-one-person-out cross-validation. We further investigate the contribution of various body locations on the recognition accuracy. Removing tags from the arms reduces accuracy by more than 10%, while removing the wrist tag only reduces accuracy by around 2%. Therefore, tag placements on the arms are more expressive for gesture recognition than on the wrist.
Index Terms:
RFID, gesture recognition, human-sensing, signal processing.
I Introduction
Radio based human sensing analyzes electromagnetic signals reflected from people or objects to infer position, gestures, vital signs, and emotional states, using technologies such as Wi-Fi [24], radar [19], RFID [48], Bluetooth [3], LoRa [27], or cellular systems [45]. Compared to vision-based methods, radio sensing has advantages, such as robustness to lighting conditions and occlusion. Although radio sensing systems can raise distinct privacy concerns due to their ability to sense through obstacles and detect fine-grained movements, they do not capture identifiable visual features. The applications of radio sensing extend to assisted living, healthcare and human-robot interaction [39].
Radio-based gesture recognition can be categorized into environmental and wearable methods. Environmental approaches (e.g. Wi-Fi, radar, etc) often rely on changes in signal propagation caused by movements within the sensing area. Reflections, distortions, or disruptions of an electromagnetic signal are interpreted to characterize gestures. Environmental methods are well suited for applications targeting ease of use. Wearable gesture recognition systems utilize body-worn sensors (e.g. backscatter tags) [2].
Environmental backscatter systems deploy tag arrays within the environment, and individuals perform gestures within the detection range of these arrays [51, 42, 23]. Wearable systems place the tags on the body [1] to enable differentiation between body parts. Wearable systems also allow individuals to evade sensing, e.g., for privacy reasons.
We explore the use of body-mounted Electronic Product Code (EPC) tags for gesture recognition and propose zero-padding and adaptive interpolation to address data loss (e.g., missing tags or samples) caused by polarization mismatch. To further improve recognition accuracy, we propose algorithmic interpolation and imputation strategies. Our contributions are
  • a gesture recognition system from body-worn backscattering tags and its validation in three environments with data from 17 subjects.
  • linear and exponential step-wise interpolation techniques as well as an imputation algorithm based on intra-class proximity and proximity-based inference.
  • data representations to capture the correlation between received signal strength (RSS) and phase values associated with different EPCs across successive timestamps, enabling efficient propagation through graph-based self-attention convolution.
II Related Work
Popular examples of environmental systems for motion recognition are Wi-Fi, radar and RFID. Particularly, the channel state information (CSI) correlates with movement relative to a receiver. [30]. Combined with deep learning, high recognition accuracies are achieved [44, 6]. A limitation of CSI-based systems is their domain dependence. To mitigate this, arranging antennas circularly around the sensing area may support orientation-agnostic recognition [29]. Furthermore, self-attention has been used to overcome domain dependence in CSI-based systems [14, 15].
In contrast, radar-based systems exploit range-doppler information, as well as spatial diversity in an antenna array to obtain information on the distance, angle and velocity of objects relative to the radar [17, 46, 40]. Systems based on in-phase and quadrature components [18], image processing over the time-frequency spectrum [28], as well as point-cloud based systems [31] have been proposed.
RFID-based systems use machine learning to interpret fluctuations in RSS and signal phase, utilizing an array of tags. [51, 42, 9]. Support vector machines, random forest, or CNN-based reasoning are prominent classification algorithms [23, 50, 37], but also Dynamic Time Warping (DTW) to group temporal patterns with a template database [53].
A common challenge for environmental systems is that signal reflections are superimposed from the complete environment and that relevant signal components first have to be isolated. Furthremore, these systems raise privacy concerns since their operation is opaque and no easy means to avoid being captured by the system exist.
We propose to integrate reflective tags into clothing. Therefore, relevant signals are easily separated from environmental reflections and avoidance of being captured is simply achieved by not wearing tags.
Body-worn sensors and deep learning have been widely used in health, human-computer interaction, and personalized medical care [26, 16, 35]. Particularly, Inertial Measurement Units (IMUs) are common for motion recognition [32]. Essential for a high recognition accuracy are the types of sensors and their placement on the human body [10, 25, 5], which can be addressed, e.g. using transfer learning [20] or transformer-based structures [38]. Common challenges are the need for frequent charging or battery replacement [43].
We propose the use of passive reflective tags which overcomes this limitation.
Attaching RFID tags to objects has been employed for human-object interactions [21, 52, 34]. By using orthogonally oriented antennas, object motion can be tracked [4]. Temporal gesture patterns can be distinguished using DTW [7, 41] and rotation agnostic recognition is possible by normalizing the horizongtal rotation angle and radial distance [49]. Coarse-grained gesture recognition utilizing machine learning and body-attached reflective tags has been presented in [47, 12].
We propose integrating reflective tags into clothing, thereby allowing the system to distinguish between various body parts using RSS and phase signals. [11, 13].
Challenges of recognition systems using body-worn tags are frequent orientation changes of the tags, which may cause poor signal quality due to polarization mismatch [8]. Furthermore, the anatomical structure and tissue composition of the human body affects signal reflection patterns [36].
In this work, we employ a graph neural network to leverage the correlations between RSS and phase values from different tags across consecutive timestamps in a gesture pattern. Furthermore, we address the problem of missing data.
Notation
Vectors and matrices are denoted by lowercase and uppercase bold letters, respectively. The $i$-th element of vector $\mathbf{a}$ is represented by $\mathbf{a}]_{i}$. For a matrix $\mathbf{A}$, $\mathbf{A}(:,i:j)$ selects all elements in all rows and columns $i$ through $j$ (inclusive), while $\mathbf{A}(i:j,:)$ selects all elements in rows $i$ through $j$ (inclusive) and all columns. The elements in the $\ell$-th row and $\ell$-th column are denoted by $\mathbf{A}(\ell,:)$ and $\mathbf{A}(:,\ell)$, respectively. For a vector $\mathbf{a}$, we use the notation $\mathbf{a}_{i:j}$ to denote the subsequence consisting of the elements from position $i$ to position $j$ (inclusive). The transpose operator is $(\cdot)^{\top}$. The $L_{2}$ norm (Euclidean norm for vectors, Frobenius norm for matrices) is denoted by $\|\cdot\|_{2}$. The cardinality of a set or the length of a vector is denoted by $|\cdot|$. The $L_{0}$ norm, $\|\mathbf{a}\|_{0}$, denotes the number of non-zero elements in $\mathbf{a}$. A column vector of length $Q$ with all elements equal to one is represented by $\mathbf{1}_{Q}$, and a column vector or matrix of size $Q$ with all elements equal to zero is represented by $\mathbf{0}_{Q}$. Assignment, where the value of $b$ replaces the value of $a$, is denoted by $a\leftarrow b$. In this paper, for two vectors $\mathbf{a}$ and $\mathbf{b}$, we define $\mathbf{a}\setminus\mathbf{b}$ as a vector containing the elements of $\mathbf{a}$ that are not present in $\mathbf{b}$, preserving their original order from $\mathbf{a}$. Note that we assume $\mathbf{b}$ is a sub-sequence of $\mathbf{a}$ for this operation. For example, if $\mathbf{a}=[1,2,3,4]$ and $\mathbf{b}=[1,4]$, then $\mathbf{a}\setminus\mathbf{b}=[2,3]$. The function $\operatorname{argsort}(\mathbf{a})$ returns the indices that would sort the vector $\mathbf{a}$ in ascending order. The complement of a set $a$ is denoted by $a^{c}$, and $\emptyset$ denotes the empty set. For two sets AA and $B$, the notation $A\subseteq B$ means that AA is a subset of $B$.
III Overview of the measurement System
We utilize Alien AZ 9662 passive RFID tags as well as an Impinj Speedway R420 RFID reader and a circularly polarized Vulcan RFID PAR90209H antenna with a circular polarization gain of 99 dBiC as well as elevation and azimuth beamwidth of $70^{\circ}$. The reader is controlled by a laptop running the Impinj ItemTest software. The RFID system operates in the 865MHz range at $30$ dBm.
The integrated circuit (IC) of the tag modulates its response to the continuous wave signal emitted by the reader by altering the load on its antenna to either backscatter or absorb the reader’s signal, thereby producing reflection coefficient $\Gamma(t)\in\{0,1\}$. The reflection coefficient is $\Gamma=\frac{Z_{\text{Ant}}-Z_{\text{Load}}(t)}{Z_{\text{Ant}}+Z_{\text{Load}}(t)}$, where $Z_{\text{Ant}}$ and $Z_{\text{Load}}(t)$ denote the antenna impedance (around $50\Omega$) and the load impedance controlled by the tag’s IC. $Z_{\text{Load}}=Z_{\text{Ant}}$ will result in $\Gamma=0$ (no signal reflected) while $Z_{\text{Load}}=0$ ($\Gamma=1$) causes the signal of the reader to be reflected. [^1] In practice, due to imperfect load matching, the maximum value of the $\Gamma$ is close to 11 and the minimum value is close to zero.
We denote the basedband received signal at the reader by
$y(t)$ $=\sum_{i=1}^{N_{p}}g_{i}G_{a}\sqrt{P_{t}}s(t)+\eta(t);\ \ \ \ y(t)\in\mathbb{C}.$ (1)
In equation (1), $g_{i}\triangleq\frac{\lambda}{4\pi d_{i}}e^{\jmath\theta_{i}}$ denotes the channel gain of the $i$th round-trip path with $\theta_{i}\sim\mathcal{U}[0,2\pi)$, wavelength $\lambda$, path distance $d_{i}$, the number of multipath components $N_{p}$, antenna gain $G_{a}$, reader transmit power $P_{t}$ and the transmitted signal $s(t)$. We model additive thermal noise at the reader as a zero-mean Gaussian $\eta(t)$ with variance $\sigma^{2}$.
The data received from the tag include EPCs, a timestamp, the signal phase, as well as the RSS.
We collected 7,140 gesture samples from 17 participants (10 males, 7 females) with varying weight and height (155cm to 185cm), each performing 21 gestures 20 times (cf. Fig. 1). Two tags were attached to the wrist of each hand, one below the elbow of each arm, and one on the upper arm of each hand as illustrated in Fig 2. Gestures were performed at 3 and 1.5 meters distance from the antenna. The duration of gestures varied due to varying gesture speed and gesture complexity.
(h) h
Figure 1: The gestures performed by participants are (a) lateral raise (1), (b) push down (2), (c) lift (3), (d) pull (4), (e) push (5), (f) lateral to front (6), (g) swipe right (7), (h) swipe left (8), (i) throw (9), (j) arm swings (10), (k) two-hand throw (11), (l) two-hand push (12), (m) two-hand pull (13), (n) two-hand lateral raise (14), (o) left-arm circle (15), (p) right-arm circle (16), (q) two-hand outward circles (17), (r) two-hand inward circles (18), (s) two-hand lateral to front (19), (t) clockwise circle (20), and (u) counterclockwise circle (21).
Figure 2: Experimental setup. Eight passive RFID tags are attached to the subject, with phase and RSS data collected using an Impinj Speedway R420 reader and a circularly polarized antenna. Participants performed gestures at distances of 1.5m and 3m from the antenna.
IV Data Preparation
To facilitate further data analysis, the raw data from the reader is processed first. Particularly, this regards, data formatting, phase unwrapping, normalizations, handling of missing samples, and interpolation. The processing framework is illustrated in Fig 3.
Initially, we assign each EPC with a numerical identifier ranging from 1 to 8 (cf. Fig 2). The RFID reader inventories tags sequentially, detecting only one EPC per timestamp. Additionally, due to movement, certain tags may become temporarily invisible during specific gestures caused by polarization mismatch between the antennas. As a result, EPCs have missing values at different timestamps. The data are initially ordered based on timestamps rather than EPCs. However, for subsequent processing tasks, it is necessary to organize the time-series data based on EPC.
Hence, the $i$-th gesture data $\mathbf{D}_{i}=[\mathbf{t}_{i},\mathbf{e}_{i},\mathbf{r}_{i},\bm{\phi}_{i}]\in\mathbb{R}^{T\times 4}$, which corresponds to $i$-th gesture, consists of RSS and phase vectors $\mathbf{r}_{i}\in\mathbb{R}^{T}$ and $\bm{\phi}_{i}\in\mathbb{R}^{T}$ for the different EPCs $\mathbf{e}_{i}\in\mathbb{R}^{T}$ at various timestamps $\mathbf{t}_{i}=[t_{1},t_{2},\dots,t_{T}]^{\top}$, where $\mathbf{t}_{i}$ is the complete sequence of timestamps from the beginning to the end of the $i$th raw gesture, and $T$ is the total number of timestamp indices.
We sort $\mathbf{D}_{i}$ based on $\text{row}_{\text{index}}=\operatorname{argsort}(\mathbf{e}_{i})$ to achieve $\tilde{\mathbf{D}}_{i}\triangleq\mathbf{D}_{i}(\text{row}_{\text{index}},:)=[\tilde{\mathbf{t}}_{i},\tilde{\mathbf{e}}_{i},\tilde{\mathbf{r}}_{i},\tilde{\bm{\phi}}_{i}]\in\mathbb{R}^{T\times 4}$. Particularly, $\tilde{\mathbf{D}}_{i}^{n}=[\tilde{\mathbf{t}}^{n}_{i},\tilde{\mathbf{e}}^{n}_{i},\tilde{\mathbf{r}}^{n}_{i},\tilde{\bm{\phi}}^{n}_{i}]\in\mathbb{R}^{M_{n}\times 4}$ is the $n$th dataframe from the $i$th gesture data, where $\tilde{\mathbf{D}}_{i}\triangleq[(\tilde{\mathbf{D}}^{1}_{i})^{\top},\dots,(\tilde{\mathbf{D}}^{N}_{i})^{\top}]^{\top}$, $M_{n}$ is the number of readings corresponding to the $n$th tag, and $\sum_{n=1}^{N}M_{n}=T$. Therefore, a single gesture data point comprises eight dataframes, where each dataframe corresponds to an EPC and includes the associated RSS and phase values as well as a timestamp.
Next, phase unwrapping is applied followed by normalization. We employ median absolute deviation (MAD) to normalize the signal phase due to its resilience to outliers:
$\beta_{t}$ $=$ $\text{median}\left(\left|[\tilde{\bm{\phi}}_{i}]_{t}-\text{median}(\tilde{\bm{\phi}}_{i})\right|\right)\quad\forall t\in\{1\dots T\}$
${[\tilde{\bm{\phi}}_{i}]_{t}}$ $=$ $\frac{[\tilde{\bm{\phi}}_{i}]_{t}-\text{median}(\tilde{\bm{\phi}}_{i})}{\beta_{t}}.$ (2)
The RSS data is normalized via Min-Max normalization:
$\tilde{\mathbf{r}}^{n}_{i}]_{m}\leftarrow\frac{[\tilde{\mathbf{r}}^{n}_{i}]_{m}-\min(\tilde{\mathbf{r}}^{n}_{i})}{\max(\tilde{\mathbf{r}}^{n}_{i})-\min(\tilde{\mathbf{r}}^{n}_{i})}\quad\forall m=\{1,2,\dots,M_{n}\}.$ (3)
Next, we employ Savitzky-Golay (S-G) [33] and Gaussian filters [22] to remove high-frequency noise:
$\tilde{\bm{\phi}}^{n}_{i}\leftarrow\mathfrak{G}(\mathfrak{S}(\tilde{\bm{\phi}}^{n}_{i})).$ (4)
In equation (4), $\mathfrak{G}(.)$ and $\mathfrak{S}(.)$ represent the Gaussian and S-G filters.
For missing RSS and phase samples, we apply zero padding:
$\bm{\varphi}^{n}_{i}(t)=\begin{cases}\tilde{\bm{\phi}}^{n}_{i}(t),&\text{if }t\in\tilde{\mathbf{t}}^{n}_{i}\\ 0,&\text{if }t\in\tilde{\mathbf{t}}_{i}\setminus\tilde{\mathbf{t}}^{n}_{i}\end{cases}$ (5)
$\bm{\rho}^{n}_{i}(t)=\begin{cases}\tilde{\mathbf{r}}^{n}_{i}(t),&\text{if }t\in\tilde{\mathbf{t}}^{n}_{i}\\ 0,&\text{if }t\in\tilde{\mathbf{t}}_{i}\setminus\tilde{\mathbf{t}}^{n}_{i},\end{cases}$ (6)
where $\bm{\rho}^{n}_{i}$ and $\bm{\varphi}^{n}_{i}$ represent the zero-padded RSS and phase data of the $n$-th dataframe of the $i$-th gesture. This results in dataframes that either contain only zero values, null dataframe, or contain a mixture of zero and non-zero values, sparse dataframe. While the null dataframe is addressed using within-class imputation methods in combination with spatial proximity imputation, the sparse dataframe is processed using linear and exponential interpolation techniques.
Figure 3: Overview of the processing framework. Hardware-induced distortions are corrected, and missing data are addressed by stepwise interpolation and spatially guided data imputation.
IV-A Interpolation
In each sparse dataframe, zero values occurring between two valid observations of phase and RSS are addressed through linear and exponential interpolation, respectively. However, in many cases, a dataframe may begin or end with a sequence of zeros, making interpolation infeasible due to the lack of two reference points. For dataframes with leading or trailing zero phase values, we apply linear extrapolation according to the first (last) two non-zero phase values, i.e.,
$\bm{\varphi}^{n}_{i}\leftarrow\mathcal{F}(\bm{\varphi}^{n}_{i})$ (7)
where $\mathcal{F}(.)$ performs the linear interpolation and extrapolation. This approach is similarly applied to RSS values, using exponential extrapolation to fill leading and trailing zero values, i.e.,
$\bm{\rho}^{n}_{i}\leftarrow\mathcal{G}(\bm{\rho}^{n}_{i})$ (8)
where $\mathcal{G}(.)$ performs the exponential interpolation and extrapolation. Finally, dataframes are resampled to $l_{\text{rs}}$.
IV-B Data Imputation
We applied within-class imputation along with spatial proximity imputation to fill the null dataframes. For each reference gesture data, we identify the null dataframes (i.e., those EPCs that are entirely missing during the inventory round) and the sparse dataframe (i.e., dataframes with available EPCs that have already undergone interpolation in the previous step).
To address data imputation, we introduce two metrics. The first metric quantifies the similarity between two gesture samples using the Mean Euclidean Distance (MED). Let $\bar{\mathbf{D}}^{a}_{i}\in\mathbb{R}^{k^{a}_{i}\times 4}$ represent a matrix that contains all sparse dataframes for the $i$-th gesture data, where the indices of their EPCs are contained in the set $\Delta_{i}$, with $|\Delta_{i}|=k^{a}_{i}$. Similarly, for the $m$-th gesture data ($m\in{1,2,\dots,M}$, $m\neq i$), with EPC indices corresponding to all sparse dataframes in the set $\Delta_{m}$, where $|\Delta_{m}|=k^{a}_{m}$ and $\Delta_{i}\subseteq\Delta_{m}$, we construct a matrix $\bar{\mathbf{D}}^{a^{\prime}}_{m}\in\mathbb{R}^{k^{a}_{i}\times 4}$ by vertically concatenating all dataframes with EPCs whose indices belong to the set $\Delta_{i}$. The MED between gesture data $i$ and $m$ is then calculated as:
$\kappa_{i,m}$ $=\frac{1}{k^{a}_{i}}\sum_{q=1}^{k^{a}_{i}}\|\bar{\mathbf{D}}^{a}_{i}(q,3:4)-\bar{\mathbf{D}}^{a^{\prime}}_{m}(q,3:4)\|_{2}.$ (9)
The second metric considers the spatial proximity of the tags to perform data imputation. To do so, we define proximity sensors based on the placement of the tags on human clothing, as illustrated in Fig. 2. As such, the proximity sensors matrix can be defined $\mathbf{A}=[\mathbf{p}_{1},\mathbf{p}_{2},\mathbf{p}_{3},\mathbf{p}_{4}]$ where $\mathbf{p}_{1}=[1,2]^{\top}$, $\mathbf{p}_{2}=[3,4]^{\top}$, $\mathbf{p}_{3}=[5,6]^{\top}$, and $\mathbf{p}_{4}=[7,8]^{\top}$. Now we consider cases where we need to perform imputation using the aforementioned metrics in the sequel.
In the first case, the $i$-th reference gesture data contains only one null dataframe for the EPC with index pp. We first calculate all MED values between the $i$-th reference gesture data and the other gesture data within the same class, i.e., $\kappa_{i,m}\quad\forall m\neq i$ given in (9). We then define the matrix $\bm{K}$, where each row contains $\kappa_{i,m}$ in the first column and the corresponding index $m$ in the second column. We identify the gesture data indices corresponding to the $\nu$ smallest values in the first column of $\bm{K}$ (i.e., $\bm{K}(:,1)$) by sorting it in ascending order. Let $v_{1}\leq v_{2}\leq\dots\leq v_{\nu}$ be these $\nu$ smallest MED values, and let $\bm{\zeta}=[\zeta_{1},\zeta_{2},\dots,\zeta_{\nu}]^{\top}$ be the vector of the corresponding gesture data indices, such that, for example, $\kappa_{i,\zeta_{1}}=v_{1}$.
Then, we define the set $\Omega\triangleq\{\forall\zeta_{j}\mid\tilde{\mathbf{D}}^{p}_{\zeta_{j}}(:,3:4)\neq\mathbf{0},j=1,2,\dots,\nu\}$, where $\tilde{\mathbf{D}}^{p}_{\zeta_{j}}(:,3:4)\neq\mathbf{0}$ indicates that the $\zeta_{j}$-th gesture in $\Omega$ contains a sparse dataframe for the EPC with index pp. Finally, the missing value $\tilde{\mathbf{D}}^{p}_{i}(:,3:4)$ is imputed by averaging the RSS and phase values from the set $\Omega$:
$\tilde{\mathbf{D}}^{p}_{i}(:,3:4)\leftarrow\frac{1}{|\Omega|}\sum_{\forall\zeta_{j}\in\Omega}\tilde{\mathbf{D}}^{p}_{\zeta_{j}}(:,3:4)$ (10)
If the MED metric cannot be computed because $\Omega=\emptyset$, the missing RSS and phase values are imputed using data from the nearest EPC within the same reference gesture, based on the proximity matrix $\mathbf{A}$.
In the second case, the reference gesture data contains two null dataframes. To this end, we first determine whether the two null dataframes belong to the same EPC pair according to $\mathbf{A}$ (i.e., whether they are spatially close based on tag placement). If the null dataframes are not from the same pair, their corresponding dataframes are then filled using the approach mentioned in the first case (see (10)), if applicable. If the MED metric cannot be applied (due to a lack of available data), the null dataframes are imputed using the dataframes from the closest available EPC within the same reference gesture data based on $\mathbf{A}$.
In contrast, when the two null dataframes belong to the same EPC pair, a scenario with low probability as indicated in Fig. 5, we initially assess the applicability of the MED metric for both. If the MED metric is applicable to both null dataframes, they are imputed using the methodology detailed in the first case (see (10)). If the MED metric is applicable to only one null dataframe (e.g., the first), it is imputed following the procedure described in the first case (see (10)). This imputed dataframe is then used to impute the second null dataframe, leveraging the spatial closeness of their EPCs. Finally, in the event that the MED metric is not applicable to either of null dataframes, the imputation is performed by averaging the values corresponding to the missing EPCs across all available gesture data within the same class. The second case can be generalized to instances that there are more than two null dataframes, on the condition that no proximity-paired EPCs are missing. Data imputation is summarized in Algorithm 2, and our model is summarized in Algorithm 1.
0:$i$-th raw gesture data $\mathbf{D}_{i}\in\mathbb{R}^{T\times 4}$, NN, and $l_{\text{rs}}$ Extract $i$-th sorted gesture data, $\tilde{\mathbf{D}}_{i}$, and dataframes $\tilde{\mathbf{D}}^{n}_{i}$ for all $n\in\{1,2,\dots,N\}$
1: Obtain the row indices that sort the EPCs $\mathbf{D}_{i}(:,2)=\mathbf{e}_{i}$ in ascending order: $\text{row}_{\text{index}}=\operatorname{argsort}(\mathbf{e}_{i})$.
2: Sort the rows of $\mathbf{D}_{i}$ based on $\text{row}_{\text{index}}$, i.e., $\tilde{\mathbf{D}}_{i}=\mathbf{D}_{i}(\text{row}_{\text{index}},:)$.
3: Define dataframes corresponding to each EPC number as: $\tilde{\mathbf{D}}_{i}\triangleq[(\tilde{\mathbf{D}}^{1}_{i})^{\top},\dots,(\tilde{\mathbf{D}}^{N}_{i})^{\top}]^{\top}$.
Phase Processing of gesture data and dataframes
4:for $n=1$ to NN do
5:  Perform phase unwrapping: $\tilde{\bm{\phi}}^{n}_{i}\leftarrow\text{unwrap}(\tilde{\bm{\phi}}^{n}_{i})$, and update $\tilde{\mathbf{D}}^{n}_{i}(:,4)\leftarrow\tilde{\bm{\phi}}^{n}_{i}$.
6:end for
7: Update $\tilde{\mathbf{D}}_{i}$ based on the updated dataframes.
8: Perform phase normalization over $\tilde{\bm{\phi}}_{i}$ based on Equation (2) and update $\tilde{\mathbf{D}}_{i}$ accordingly.
9:for $n=1$ to NN do
10:  Perform phase smoothing based on Equation (4) and update $\tilde{\mathbf{D}}^{n}_{i}(:,4)\leftarrow\tilde{\bm{\phi}}^{n}_{i}$, and $\tilde{\mathbf{D}}_{i}$.
11:end for
RSS Normalization
12:for $n=1$ to NN do
13:  Perform RSS normalization based on Equation (3), update $\tilde{\mathbf{D}}^{n}_{i}(:,3)\leftarrow\tilde{\mathbf{r}}^{n}_{i}$, and update $\tilde{\mathbf{D}}_{i}$.
14:end for
Zero Padding $\tilde{\mathbf{D}}^{n}_{i}$ for $n\in\{1,2,\dots,N\}$
15:for $n=1$ to NN do
16:  Perform zero-padding using Equations (5) and (6), resulting in $\tilde{\mathbf{D}}^{n}_{i}\leftarrow[\tilde{\mathbf{t}}_{i},n\mathbf{1}_{T},\bm{\rho}^{n}_{i},\bm{\varphi}^{n}_{i}]\in\mathbb{R}^{T\times 4}$.
17:  Perform zero padding validation:
18:  if $\|\bm{\rho}^{n}_{i}\|_{0}<2$ or $\|\bm{\varphi}^{n}_{i}\|_{0}<2$ then
19:     $\bm{\rho}^{n}_{i}\leftarrow\mathbf{0}_{T}$ and $\bm{\varphi}^{n}_{i}\leftarrow\mathbf{0}_{T}$
20:  end if
21:end for
Perform interpolation/extrapolation and data imputation on $\tilde{\mathbf{D}}^{n}_{i}$ for all $n\in\{1,2,\dots,N\}$
22:for $n=1$ to NN do
23:  Perform interpolation/extrapolation on $\bm{\rho}^{n}_{i}$ and $\bm{\varphi}^{n}_{i}$.
24:  if $\|\bm{\rho}^{n}_{i}\|_{0}>0$ or $\|\bm{\varphi}^{n}_{i}\|_{0}>0$ then
25:      Perform Equations (7) and (8), and update $\tilde{\mathbf{D}}^{n}_{i}$ and $\tilde{\mathbf{D}}_{i}$.
26:   $\bm{\varphi}^{n}_{i}\leftarrow\text{Res}(\bm{\varphi}^{n}_{i},l_{\text{rs}})$ and $\bm{\rho}^{n}_{i}\leftarrow\text{Res}(\bm{\rho}^{n}_{i},l_{\text{rs}})$, $\tilde{\mathbf{D}}^{n}_{i}(:,3:4)\leftarrow[\bm{\rho}^{n}_{i},\bm{\varphi}^{n}_{i}]$, and update $\tilde{\mathbf{D}}_{i}$.
27:  else
28:      Skip interpolation/extrapolation and proceed to data imputation (Algorithm 2).
29:  end if
30:end for
Return: Refined processed $i$-th gesture data $\tilde{\mathbf{D}}_{i}$.
Algorithm 1 the Proposed Model
0: All null dataframes $\tilde{\mathbf{D}}^{n}_{i}$ where $n\in\Delta^{c}_{i}$ (set of missed EPC numbers, $|\Delta^{c}_{i}|=\epsilon_{i}$), all gesture data sets in the training phase, available EPC numbers in the $i$-th gesture data: $\Delta_{i}$, Proximity sensor matrix $\mathbf{A}$, and the number of nearest gesture data to consider: $\nu$.
1: Extract the matrix of sparse dataframes for the $i$-th and $m$-th gesture data: $\bar{\mathbf{D}}^{a}_{i},\bar{\mathbf{D}}^{a^{\prime}}_{m}\in\mathbb{R}^{k^{a}_{i}\times 4}$, where $\Delta_{i}\subseteq\Delta_{m}$.
2: Calculate $\kappa_{i,m}$ for all $i,m\in\{1,2,\dots,M\}$, and $m\neq i$, based on Equation (9). Store $\kappa_{i,m}$ and $m$ in the first and the second column of matrix $\bm{K}$, respectively.
3: Find the $\nu$ row indices corresponding to the $\nu$ smallest values, $\mathcal{I}_{\nu}\triangleq[\operatorname{argsort}(\bm{K}(:,1))]_{1:\nu}$, and store the indices in: $\bm{\zeta}\triangleq\bm{K}(\mathcal{I}_{\nu},2)$
4: Define the set of nearest neighbor gesture data indices $\Omega\triangleq\{\forall\zeta_{j}\mid j=1,2,\dots,\nu\}$, where $\zeta_{j}\triangleq[\bm{\zeta}]_{j}$.
5:for $\forall\mu\in\Delta^{c}_{i}$ do
6:  Initialize the set of candidate gesture data for imputation: $\Omega_{\mu}\leftarrow\Omega$.
7:  Update $\Omega_{\mu}\leftarrow\{\forall\zeta_{j}\in\Omega\mid\tilde{\mathbf{D}}^{\mu}_{\zeta_{j}}(:,3:4)\neq\mathbf{0}\}$.
8:  if $\Omega_{\mu}\neq\emptyset$ then
9:   Impute the null dataframe for the $\mu$-th EPC in the $i$-th gesture data $\tilde{\mathbf{D}}^{\mu}_{i}(:,3:4)$ by averaging over the data in $\Omega_{\mu}$ according to Equation (10), and update $\tilde{\mathbf{D}}_{i}$.
10:  else
11:   if $\tilde{\mathbf{D}}^{\delta}_{i}\neq\mathbf{0}$, where $\delta$ and $\mu$ belong to the same column of matrix $\mathbf{A}$ then
12:    Impute: $\tilde{\mathbf{D}}^{\mu}_{i}\leftarrow\tilde{\mathbf{D}}^{\delta}_{i}$, and update $\tilde{\mathbf{D}}_{i}$.
13:   else
14:    Perform Equation (10) over all gesture data that have available $\mu$-th EPC, and update $\tilde{\mathbf{D}}_{i}$.
15:   end if
16:  end if
17:end for
Return: Refined version of the $i$-th dataframe $\tilde{\mathbf{D}}_{i}$.
Note 1: Algorithm 2 is applicable when the number of missing EPCs $\epsilon_{i}\leq 4$ and these missing EPCs are not proximity tags.
Note 2: Line 14 is applicable only during training, where this scenario’s probability is low (see Fig. 5). During testing, only Line 12 is used, which is practical (see Fig. 5).
Algorithm 2 Null Dataframe Imputation
V Attention-Based Graph Learning
For the classification, we employ the graph-based convolutional neural network introduced in [31]. We use temporal K-nearest neighbors (K-NN)  [^2] We use Euclidean distance as the metric to determine neighbors in the K-NN. to form a graph where each EPC (tag) is represented as a node, and edges are established between nodes based on the proximity of their RSS and phase values across consecutive timestamps [^3] In this study, with a large temporal hyperparameter, K-NN is applied to find neighbors between consecutive timestamps. Specifically, for two EPCs $i$ and $j$, an edge is formed if the signal feature of EPC $i$ at time t1t-1, $(\rho_{i}^{t-1},\varphi_{i}^{t-1})$, is among the $K$ nearest neighbors of the signal feature of EPC $j$ at time $t$, $(\rho_{j}^{t},\varphi_{j}^{t})$, in the feature space.. As illustrated in Fig. 4, the framework connects each EPC to its most similar counterparts at the next timestamp, thereby capturing meaningful temporal correlations and enabling efficient feature propagation across the graph.
Afterward, it employs graph processing including message propagation, self-attention, and an aggregation function. In the message propagation step, node $i$ receives information from node $j$, where the information is a function of the features of these nodes. This function is implemented by a multilayer perceptron. To focus on the most relevant neighbors, a self-attention mechanism is incorporated. This mechanism computes attention weights for each neighbor $j$ of node $i$, scaling the influence of the corresponding message. The updated feature vector for node $i$ at the next layer is then obtained by aggregating the weighted messages and the node’s previous feature [^4] The aggregation function can be mean or summation..
To facilitate model training, the input data is structured into a four-dimensional tensor $\bm{\mathcal{T}}\in\mathbb{R}^{B\times T\times N\times D}$ where $B$ denotes the total number of gesture samples, $T$ is the number of timestamps per EPC in each gesture sample, NN represents the number of EPCs, and $D$ is the feature dimension of each EPC (i.e., RSS and phase). In this work, the tensor shape used is ($T=30$, $N=8$, $D=2$), corresponding to 30 timestamps, 8 EPCs, and two-dimensional feature vectors per node.
Figure 4: Graph generation based on $(\rho_{i},\varphi_{i})$ values across consecutive timestamps. For each EPC at a given timestamp, directed edges are formed by connecting to the nearest neighbors in the preceding timestamp using a temporal graph K-NN approach.
VI Evaluation
We present accuracy, precision, recall, and F1-score for within-subject training in Table I(a), Table I(b), and Table I(c). The confusion matrix for the single-handed gestures (Fig 6(a)) demonstrates a within-user test accuracy of 98.27%. Notably, perfect classification is achieved for seven gestures, while the remaining classes accuracies exceed 90%. The confusion matrix for the complete set of gestures (Fig 6(c)) yields an overall within-user test accuracy of 98.13%. Out of 21 gestures, 16 were classified with 100% accuracy, and rest with accuracies greater than 90%.
The performance is evaluated by the Random Forest Classifier models from [50]. Particularly, we employed a classifier utilizing exclusively statistical features derived from the phase signal (RFC with SP), both statistical features and wavelet coefficients obtained from the phase (RFC with SWP), as well as statistical featues from both phase and RSS (RFC with SPR). In addition, the model is evaluated using Early Fusion [5], Late Fusion [13], EUIGR [47], and GRfid [53]. All models were implemented and evaluated across three distinct datasets, constructed as follows:
Dataset I: Gesture data were collected at a distance of 3 meters in Environment A. Dataset II: Gesture data were collected at a distance of 1.5 meters in Environment A. Dataset III: Gesture data were collected at a distance of 1.5 meters in Environment B.
Our model consistently outperforms the other models across all datasets (cf. Table I(a), Table I(b), and Table I(c)). Among the random forest models, RFC with SPR, including more features and additional modalities improves the recognition accuracy. Late Fusion outperforms Early Fusion while EUIGR and GRfid exhibit notably lower performance, with GRfid in particular performing poorly across all metrics (e.g., 30.34% accuracy).
The confusion matrices in Fig 6(b) and Fig 6(d) depict the performance of our model following leave-one-person-out cross-validation on Dataset I for single-handed and complete set of gestures, achieving test accuracies of 94.44% and 89.28%.
VI-A Parameter tuning
Fig 7 shows the model’s performance for different configurations of the top-$\nu$ nearest data points and the resampling length, $l_{\text{rs}}$. For datasets I and III, the highest accuracies were achieved with $l_{\text{rs}}=30$ (98.13% and 98.41%). For Dataset II, the best performance was obtained with $l_{\text{rs}}=20$ (96.82%). Overall, $l_{\text{rs}}$ values of 20 and 30 demonstrated superior performance compared to the settings of 10 and 40.
Additionally, increasing $\nu$ generally improved accuracy, with the optimal performance observed at $\nu=30$.
VI-B Impact of tag placement and misdetection
Fig 5 depicts the proportion of gesture instances in which one or more tags were not detected. Among all tags, tag 6 exhibited the highest rate of loss (42%), followed by tag 2 (15%) due to the placement of these tags at the back of the right and left wrist (cf. Fig 2). Missed detections were also observed for Tags 1 and 5, located on the front side of the wrists, with loss rates of 9% and 15%, respectively. In contrast, Tags 3, 4, 7, and 8, which are positioned on the forearms and upper arms, demonstrated significantly lower loss rates.
Notably, as shown in Fig 5, instances where any pair of tags is lacking data at the same time are rare, so that indeed imputation is well feasible drawing on the samples collected by other tags.
Fig 8 evaluates how the system’s accuracy changes across all datasets when specific tags or combinations of tags are omitted. It is evident that certain tags contribute more to the overall performance of the recognition system. Notably, T4 and T8 consistently exhibit the highest individual impact on accuracy. For instance, in Dataset I (Fig 8(a)), omitting either T4 or T8 leads to a noticeable drop in accuracy from approximately 97.8% to around 95%, and further to  94% when both are removed. Following T4 and T8, T3 and T7 also demonstrate relatively higher importance compared to tags placed on the wrist, with their combined removal resulting in an accuracy of approximately 94.6%. Overall, in all datasets, the ”No Tag” condition, where none of the tags are removed, yields the highest accuracy. This indicates that utilizing all eight tags attached to various positions on the garment is beneficial for the recognition system. However, as illustrated in Fig 5, tag misdetections may occur, leading to intermittent missing values or complete detection failures at certain timestamps. These findings show the importance of our model in robustly handling missing tag data.
Figure 5: Probability of individual tag misdetection and joint tags’ misdetection across all gesture executions.
TABLE I: Accuracy (Acc.), Precision (Pre.), Recall (Rec.), and F1-score (F1) for all methods
(a) Distance $3~\text{m}$; environment A
Method Acc. Pre. Rec. F1
RFC with SP $83.56$ $83.72$ $83.56$ $83.42$
RFC with SWP $86.26$ $86.48$ $86.2$ $86.07$
RFC with SPR $95.25$ $95.35$ $95.25$ $95.23$
Early Fusion $83.62$ $84.25$ $83.57$ $83.34$
Late Fusion $87.13$ $88.32$ $87.08$ $86.91$
EUIGR $80.4$ $78.94$ $80.07$ $78.73$
GRfid $30.34$ $31.4$ $30.16$ $30.22$
Our model $\mathbf{98.13}$ $\mathbf{98.19}$ $\mathbf{98.13}$ $\mathbf{98.13}$
(b) Distance $1.5~\text{m}$; environment A
Method Acc. Pre. Rec. F1
RFC with SP $85.07$ $86.29$ $85.02$ $85.13$
RFC with SWP $85.71$ $86.72$ $85.71$ $85.72$
RFC with SPR $93.52$ $94.35$ $93.52$ $93.59$
Early Fusion $81.01$ $86.08$ $81$ $81.18$
Late Fusion $89.41$ $90.18$ $89.39$ $89.39$
EUIGR $80.37$ $74.75$ $80.33$ $75.97$
GRfid $29.39$ $29.14$ $29.34$ $28.89$
Our model $\mathbf{96.82}$ $\mathbf{97.88}$ $\mathbf{96.8}$ $\mathbf{97.02}$
(c) Distance $1.5~\text{m}$; environment B
Method Acc. Pre. Rec. F1
RFC with SP $84.12$ $84.67$ $84.07$ $83.96$
RFC with SWP $81.26$ $82.39$ $81.26$ $80.93$
RFC with SPR $93.80$ $94.08$ $93.78$ $93.71$
Early Fusion $91.13$ $94.16$ $91.13$ $91.56$
Late Fusion $90.15$ $91.13$ $90.15$ $90.15$
EUIGR $87.30$ $83.75$ $87.27$ $84.55$
GRfid $34.43$ $36.3$ $33.87$ $34.33$
Our model $\mathbf{98.41}$ $\mathbf{98.67}$ $\mathbf{98.41}$ $\mathbf{98.39}$
(a) Within-subject; Accuracy=98.27%
(b) Leave-one-person-out; Acc=94.44%
(c) Within-subject; Accuracy=98.13%
(d) Leave-one-person-out; Acc=89.28%
Figure 6: Normalized confusion matrices for the single-handed and complete sets.
(a) 3m distance to the reader
(b) 1.5m distance to the reader
(c) 1.5m distance in a different environment
Figure 7: Test accuracy results for varying combinations of $\nu$ and $l_{\text{rs}}$
(a) 3m distance to the reader
(b) 1.5m distance to the reader
(c) 1.5m distance in a different environment
Figure 8: Test accuracy under the removal of individual tags and sets of tags.
VII Conclusion
We have explored the use of body-mounted UHF RFID tags for human gesture recognition. Leveraging a commercial RFID reader and multiple passive tags positioned on the body, we recognized 21 distinct gestures. We proposed a processing framework to handle partially missing dataframes using linear and exponential interpolation/extrapolation for phase and RSS values. Fully missing frames were addressed through imputation and proximity-based inference. The processed data were structured as graph, with each EPC treated as a node and edges formed based on temporal correlations between RSS and phase values across successive timestamps. This graph representation was then utilized by a self-attention-based graph convolutional network. We achieved a recognition accuracy of 98.13%.